# Investigating and Mitigating Stereotype-aware Unfairness in LLM-based Recommendations

### v1

@@ Abstract.
Large Language Models (LLMs) have demonstrated unprecedented language understanding and reasoning capabilities to capture diverse user preferences and advance personalized recommendations. Despite the growing interest in LLM-based personalized recommendations, unique challenges are brought to the trustworthiness of LLM-based recommender systems (LLM-RS), since LLMs are likely to inherit stereotypes that are embedded ubiquitously in word embeddings due to their training on large-scale uncurated datasets. This leads to LLM-RS exhibiting stereotypical linguistic associations between users and items. However, there remains a lack of studies investigating the simultaneous existence of stereotypes in the word embeddings of user and item in LLM-RS. To bridge this gap, this study reveals a new variant of fairness between stereotype groups containing both users and items, to quantify discrimination against stereotypes in LLM-RS. Moreover, in this paper, to mitigate stereotype-aware unfairness in textual user and item information, we propose a novel framework (MoS), in which an insightful stereotype-wise routing strategy over multiple stereotype-relevant experts is designed to learn unbiased representations against different stereotypes in LLM-RS. Extensive experiments are conducted to analyze the influence of stereotype-aware fairness in LLM-RS and the effectiveness of our proposed methods, which consistently outperform competitive benchmarks under various fairness settings.
@@
Recommendation, Large Language Model, Fairness, Stereotype, Recommender System.
†Correspondence to: Wenqi Fan and Qing Li, Department of Computing, The Hong Kong Polytechnic University.
copyright: acmlicensedjournalyear: 2018isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. INTRODUCTION
Figure 1. Illustration of stereotype-aware fairness. When a user’s historical interactions tend to be grouped in the female stereotype due to the dominance of stereotypical textual knowledge, LLM-RS tend to overlook the user’s personalized preferences by over-recommending items (i.e., increase from 70% to 95%) from female stereotype group and under-recommend items (i.e., decrease from 30% to 5%) from male stereotype group.
@@ 1. INTRODUCTION
Recommender systems (RS) provide personalized suggestions tailored to user preferences, facilitating user experience across diverse applications (Fan et al., 2019; Jin et al., 2024; Wu et al., 2024; Fan et al., 2020), such as e-commerce, job matching, and social media platforms. Recently, Large Language Models (LLMs) have emerged as a prevalent paradigm for advancing personalized recommendations. To be specific, LLMs equipped with billion-scale parameters have demonstrated unprecedented language understanding and reasoning capabilities to capture diverse user preferences based on rich textual side information in RS (e.g., user profiles and item descriptions) (Zhao et al., 2024; Qu et al., 2024). However, the integration of LLMs into recommendations brings about unique challenges toward the trustworthiness of LLM-based recommender systems (LLM-RS), as recent studies have revealed that LLMs trained on large-scale uncurated data inherit stereotypes against social groups, leading to intrinsic biases in downstream applications (Navigli et al., 2023; Gallegos et al., 2024). For instance, LLMs tend to overlook the personalized preference behind user-item interactions but simply perform recommendations based on stereotypical textual knowledge, such as suggesting ”female nurse” and ”male doctor” in job recommendations (Kotek et al., 2023; Gallegos et al., 2024).
@@
The fairness in recommendations can be categorized into three types according to the stakeholders involved in modeling user-item interactions, namely user-side, item-side, and two-sided fairness (Wang et al., 2023; Chen et al., 2023). Most existing studies on LLM-RS fairness focus on either user-side fairness to achieve consistent recommendation performance across user groups (Hua et al., 2023a; Zhang et al., 2023; Deldjoo and Di Noia, 2024) or item-side fairness by providing fair exposure opportunities across item groups (Jiang et al., 2024; Bao et al., 2024). However, the LLM-encoded stereotypes introduce inherent biases that can simultaneously affect user groups and item groups. To be specific, stereotypes are embedded ubiquitously in word embeddings of LLM-RS (e.g., user profiles and item titles), exhibiting stereotypical linguistic associations between users and items. As illustrated in Figure 1, we take the female stereotype group as an example. When a user’s historical interactions tend to be grouped in the female stereotype due to the dominance of stereotypical textual knowledge, LLM-RS tend to overlook the user’s personalized preferences by over-recommending items (i.e., increase from 70% to 95%) from female stereotype group and under-recommend items (i.e., decrease from 30% to 5%) from male stereotype group. This leads to the reduction in user satisfaction and diversity of personalized recommendations, since stereotypes might hinder LLM-RS from exploring potential items for users. Moreover, harmful stereotypes of LLMs could further reinforce social polarization in recommendations, such as suggesting low-paid jobs to certain gender identities and nationalities (Wang et al., 2024). Therefore, it is imperative to delve into the fairness against stereotypes in the LLM-based recommender systems.
@@
Due to the simultaneous existence of stereotypes in the word embeddings of user and item in LLM-RS, as compared in Figure 1, existing user-side or item-side fairness could fall short of modeling textual stereotypes. Therefore, we propose a new variant of fairness between stereotype groups containing both users and items (i.e., two-sided groups), rather than separating user and item groups. To validate the existence of stereotypes and quantify the recommendation unfairness between different stereotype groups, we design a new evaluation metric named stereotype-aware fairness and conduct preliminary experiments on real-world recommendation datasets. As detailed in Section 2.3, our findings demonstrate that LLM-RS can exhibit significant discrimination between different stereotype groups, highlighting the concern of stereotype-aware fairness toward the trustworthiness of LLM-RS.
@@
To mitigate unfairness caused by stereotypes, unique challenges are posed to LLM-RS, since users and items are not consolidated into user and item embeddings but a sequence of tokens of textual descriptions (e.g., item titles) (Hua et al., 2023b; Qu et al., 2024). For example, most existing fairness criteria in RS, which are used to measure the similarity between user and item embeddings, are inapplicable to token sequences (Wan et al., 2020; Wang et al., 2024). This leads to significant difficulties in distinguishing different stereotype groups based on the textual information of users and items in LLM-RS. Recently, studies have revealed that LLMs possess virtual personalities that are sensitive to prompt biases (Rao et al., 2023; Xu et al., 2023). For example, LLM agents can exhibit diverse human-like personalities by giving user profiles in prompts (Frisch and Giulianelli, 2024; La Cava et al., 2024), uncovering the great potential to assign different stereotype roles to LLMs. Building up these insights, we propose a novel method named Mixture-of-Stereotypes (MoS) to capture and mitigate different stereotypes in LLM-RS, utilizing a set of stereotype-relevant experts (i.e., multiple stereotype roles). More specifically, we develop an insightful routing strategy over multiple stereotype-relevant experts to learn unbiased representations against different stereotypes into soft prompts and integrate with LLM-RS via prompt tuning, as existing research has demonstrated the effectiveness of adapting multiple experts to the training of LLMs with parameter-efficient fine-tuning (PEFT) paradigms.
@@
  • This study investigates the unique characteristics of stereotypes in LLM-RS that simultaneously exist in the word embeddings of users and items. In this paper, we propose a new variant of fairness between stereotype groups containing both users and items (i.e., two-sided groups) in LLM-RS.
  • We propose a novel framework (MoS) to mitigate discrimination against stereotypes in LLM-based recommendations, where an insightful stereotype-wise routing strategy over multiple stereotype-relevant experts is designed to learn unbiased representations against different stereotypes in LLM-RS.
@@
2. PRELIMINARY
@@
In this section, we first elaborate on the proposed stereotype-aware fairness for quantifying discrimination against stereotypes in LLM-RS. Subsequently, a preliminary experiment is initiated to address the following two research questions:
@@
  • RQ1: Does LLM-RS exhibit unfairness between the same ($u,\hat{v}\in G$) and different ($u\in G,\hat{v}\notin G$) stereotype groups?
@@
Table 1. Results of preliminary experiment. In the preliminary setup, LLM-RS are LoRA fine-tuned (Hu et al., [n. d.]) with recommendation datasets. We report the average results over three independent runs with random seeds. The reported results are multiplied by 100, where boldface indicates the best score.
$G\in\mathcal{G}$ users $u$ target items $\hat{v}$ implicit explicit counterfactual AUC Precision Recall
MovieLens male/female* $u\in G$ $\hat{v}\notin G$4.53 -13.39 0.15 49.54 60.94 82.86
$u\in G$ $\hat{v}\in G$ 80.71 81.38 77.84 71.02 75.73 87.01
BookCrossing teen/adult* $u\in G$ $\hat{v}\notin G$ 17.62 21.36 17.62 76.38 69.26 86.52
$u\in G$ $\hat{v}\in G$ 58.67 64.57 60.00 93.75 100.00 81.25
  • * As detailed in Section 4.1, we utilize the gender and age features in MovieLens and BookCrossing datasets to divide users and items into stereotype groups, respectively.
2.1. Stereotype Group in Recommendations
2.1.1. Recommendation Task.
To adapt generative LLMs to recommendation tasks, recent advances have demonstrated the necessity to present target item candidates into prompts (Bao et al., 2023; Liao et al., 2023) or additional tokens (Zhu et al., 2024; Qu et al., 2024) of LLM-RS. Therefore, we formulate the recommendation task as a binary classification problem, where LLM-RS will determine whether or not to recommend a given target item $v$ based on the sequence of a user $u$’s historical interactions $\mathcal{H}_{u}$.
2.1.2. Stereotype Group of User and Item.
Different from previous fairness that separately considers user groups and item groups, we advance the concept of stereotype groups in LLM-RS that encompasses both users and items, since LLM-encoded stereotypes are simultaneously embedded in the word embeddings of users and items. For example, a ”female” user and an item ”lipstick” can be in the same gender stereotype group.
Formally, let $G\in\mathcal{G}$ denote each stereotype group in a recommendation dataset, the user-side stereotype that $u\in G$ can be directly determined by user attributes, such as gender and age. As for item-side stereotype, we interpret it as a degree $d_{v\in G}$ to which this item is mostly interacted by users of a certain stereotype group. For example, in the case of binary stereotype groups $\mathcal{G}=[G_{1},G_{2}]$ where an item $v$ is interacted by $30\%$ users $u\in G_{1}$ and $10\%$ users $u\in G_{2}$, the degree $d_{v\in G_{1}}=0.2$ can be calculated by the subtraction. Formally, for any stereotype groups $G,G^{\prime}\in\mathcal{G}$ ($G\neq G^{\prime}$), the degree to which an item $v$ is biased to a certain stereotype group $G$ can be calculated by
(1) $d_{v\in G}=\mathop{\max}_{G}(\frac{\sum_{u\in G}{\mathbb{1}(v\in\mathcal{H}_{u})}}{\sum_{u\in G}1}-\sum_{G^{\prime}\in\mathcal{G}}{\frac{\sum_{u^{\prime}\in G^{\prime}}{\mathbb{1}(v\in\mathcal{H}_{u^{\prime}})}}{\sum_{u\in G^{\prime}}1}}),$
2.2. Evaluation of Stereotype-aware Fairness
2.2.1. Stereotype Measurement.
Building upon the above definition of stereotype groups, the stereotype of LLM-RS can be quantified by amplifying the original preference of a user towards its own stereotype group. Formally, for any specific stereotype group $G$ that a user belongs to (i.e., $u\in G$), we measure this user’s original preference by the proportion of interacted item $v$ from the same stereotype group $G$ in the interaction history $\mathcal{H}_{u}$ as follows:
(2) $h_{u\in G}=\frac{1}{|\mathcal{H}_{u}|}\sum_{v\in\mathcal{H}_{u}}{\mathbb{1}(d_{v\in G})},$
where $\mathbb{1}$ represents an identity function
(3) $\mathbb{1}(d_{v\in G})=\begin{cases}1,&\text{if}\ d_{v\in G}\geq\mathrm{threshold}\\ 0,&\text{otherwise}\end{cases}.$
In practice, a threshold should be applied to the degree of item-side stereotype. This is because a small degree $d_{v\in G}$ indicates that this item $v$ is weakly biased to any specific stereotype group as discussed in Section 2.1.2. To determine the proper threshold, we employ the Z-score of $d_{v\in G}$ over all items in a recommendation dataset. Specifically, Z-score identifies outliers based on how many standard deviations a data point is from the mean, where we regard outliers as items strongly related to certain stereotypes.
Figure 2. Threshold of $d_{v\in G}$ based on Z-scores ($z=3$) in different experimental datasets. In Figure 5, an ablation study is conducted to validate the above threshold of stereotype measurement based on Z-scores.
As illustrated in Figure 2, we take the commonly-used setting $z=3$ (Vuyyala et al., 2023) to determine the threshold in Eq. (3) tailored to different experimental datasets. To validate the threshold of stereotype measurement based on Z-scores, we conduct an ablation study as detailed in Section 4.3.3.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated unprecedented language understanding and reasoning capabilities to capture diverse user preferences and advance personalized recommendations. Despite the growing interest in LLM-based recommendations, unique challenges are brought to the trustworthiness of LLM-based recommender systems (LLM-RS). Compared to unique user/item representations in conventional recommender systems, users and items share the textual representation (e.g., word embeddings) in LLM-based recommendations. Recent studies have revealed that LLMs are likely to inherit stereotypes that are embedded ubiquitously in word embeddings, due to their training on large-scale uncurated datasets. This leads to LLM-RS exhibiting stereotypical linguistic associations between users and items, causing a form of two-sided (i.e., user-to-item) recommendation fairness. However, there remains a lack of studies investigating the unfairness of LLM-RS due to intrinsic stereotypes, which can simultaneously involve user and item groups. To bridge this gap, this study reveals a new variant of fairness between stereotype groups containing both users and items, to quantify discrimination against stereotypes in LLM-RS. Moreover, in this paper, to mitigate stereotype-aware unfairness in textual user and item representations, we propose a novel framework named Mixture-of-Stereotypes (MoS). In particular, an insightful stereotype-wise routing strategy over multiple stereotype-relevant experts is designed, aiming to learn unbiased representations against different stereotypes in LLM-RS. Extensive experiments are conducted to analyze the influence of stereotype-aware fairness in LLM-RS and the effectiveness of our proposed methods, which consistently outperform competitive benchmarks under various fairness settings.
Index Terms:
@@
Recommender System, Large Language Model, Recommendation, Fairness, Stereotype.
I INTRODUCTION
Figure 1: Illustration of stereotype groups. Compared to unique user/item representations in conventional recommender systems, users and items share the textual representation (e.g., word embeddings) in LLM-based recommendations. Therefore, LLM stereotypes can simultaneously involve user and item groups, leading to a form of two-sided recommendation fairness.
@@ I INTRODUCTION
RECOMMENDER systems (RS) provide personalized suggestions tailored to user preferences, facilitating user experience across diverse applications, such as e-commerce [1], job matching [2], and social media platforms [3, 4]. Recently, Large Language Models (LLMs) have emerged as a prevalent paradigm for advancing personalized recommendations. To be specific, LLMs equipped with billion-scale parameters have demonstrated unprecedented language understanding and reasoning capabilities to capture diverse user preferences based on rich textual side information in RS (e.g., user profiles and item descriptions) [5]. However, the integration of LLMs into recommendations brings about unique challenges toward the trustworthiness of LLM-based recommender systems (LLM-RS). Specifically, recent studies have revealed that LLMs trained on large-scale uncurated data inherit stereotypes against social groups, leading to intrinsic biases in downstream applications [6, 7]. For instance, LLMs tend to overlook the personalized preference behind user-item interactions but simply perform recommendations based on stereotypical textual knowledge, such as suggesting “female nurse” and “male doctor” in job recommendations [7, 8].
Figure 2: Illustration of stereotype-aware fairness. In A and B, existing user-side or item-side fairness still ignores two-sided (i.e., user-to-item) stereotypes, only considering the fairness against single-sided stereotypes. In C, we model the fairness against two-sided (i.e., user-to-item) stereotypes based on calibrated proportions. For example, the user-side stereotype can unfairly amplify the original distribution of item-side stereotypes, by over-recommending items (i.e., increase from 70% to 95%) in the female stereotype group and under-recommending items (i.e., decrease from 30% to 5%) in the male stereotype group.
@@
The fairness in recommendations can be categorized into three types according to the stakeholders involved in modeling user-item interactions, namely user-side, item-side, and two-sided fairness [9, 10]. Most existing studies on LLM-RS fairness focus on either user-side fairness to achieve consistent recommendation performance across user groups [11, 12, 13] or item-side fairness by providing fair exposure opportunities across item groups [14, 15]. However, as shown in Figure 1, LLM-encoded stereotypes can simultaneously involve user and item groups, leading to a form of two-sided (i.e., user-to-item) recommendation fairness. This is because stereotypes are embedded ubiquitously in the word embeddings of LLMs [6, 7], where users and items share these representations (e.g., word embeddings) in LLM-based recommendations. For example, the same vocabulary of word embeddings is used by LLMs to represent user profiles and item descriptions. This leads to LLM-RS exhibiting unfair recommendations due to stereotypical linguistic associations between users and items, as illustrated in Figure 2, we take the female stereotype as an example. Due to stereotypes in word embeddings, LLM-RS tend to group a user’s historical interactions into the female stereotype. As a result, the user-side stereotype can unfairly amplify the original distribution of item-side stereotypes, by over-recommending items (i.e., increase from 70% to 95%) in the female stereotype group and under-recommending items (i.e., decrease from 30% to 5%) in the male stereotype group.
Failing to address the unfairness due to stereotypes in LLM-based recommendations can significantly reduce user satisfaction and diversity of personalized recommendations, since stereotypes might hinder LLM-RS from exploring potential items for users. Moreover, harmful stereotypes of LLMs could further reinforce social polarization in recommendations, such as suggesting low-paid jobs to certain gender identities and nationalities [16]. Therefore, it is imperative to delve into the fairness against stereotypes in LLM-based recommendations.
@@
Due to the simultaneous existence of stereotypes in user and item groups, as explained in Fig. 1, existing user-side or item-side fairness could fall short of modeling textual stereotypes. Therefore, we propose a new variant of fairness between stereotype groups containing both users and items (i.e., two-sided groups), rather than separating user groups and item groups. To validate the existence of stereotypes and quantify the recommendation unfairness between different stereotype groups, we design a new evaluation metric named stereotype-aware fairness and conduct preliminary experiments on real-world recommendation datasets. As detailed in Section V, our findings demonstrate that LLM-RS can exhibit significant discrimination between different stereotype groups, highlighting the concern of stereotype-aware fairness toward the trustworthiness of LLM-based recommendations.
@@
To mitigate unfairness caused by stereotypes, unique challenges are brought to LLM-based recommendations. Compared to conventional recommender systems, users and items are not consolidated into unique representations, but span over a sequence of textual representations (e.g., item descriptions) in LLM-RS [17, 18]. For example, most existing fairness criteria in RS, which are used to measure the similarity between user and item embeddings, are inapplicable to word sequences [16, 19]. This leads to significant difficulties in distinguishing different stereotype groups based on the textual information of users and items in LLM-RS. Recently, studies have revealed that LLMs possess virtual personalities that are sensitive to prompt biases [20, 21]. For example, LLM agents can exhibit diverse human-like personalities by giving user profiles in prompts [22, 23], uncovering the great potential to assign different stereotype roles to LLMs. Building up these insights, we propose a novel method named Mixture-of-Stereotypes (MoS) to capture and mitigate different stereotypes in LLM-RS, utilizing a set of stereotype-relevant experts (i.e., multiple stereotype roles). More specifically, we develop an insightful routing strategy over multiple stereotype-relevant experts to learn unbiased representations, which can be seamlessly integrated with LLMs via parameter-efficient fine-tuning (PEFT) like prompt tuning, as existing research has demonstrated the effectiveness of adapting multi-expert structures to the training of LLMs via PEFT paradigms [24, 25].
@@
  • This study investigates the unique characteristics of LLM stereotypes in recommendation unfairness, caused by the stereotypical linguistic associations between users and items. In this paper, we propose a new variant of fairness against stereotypes, which simultaneously involves user and item groups (i.e., two-sided groups) in LLM-RS.
  • We propose a novel framework, namely Mixture-of-Stereotypes (MoS), to mitigate discrimination against stereotypes in LLM-based recommendations. Notably, an insightful stereotype-wise routing strategy over multiple stereotype-relevant experts is designed to learn unbiased representations against different stereotypes in LLM-RS.
@@
II RELATED WORK
In this section, we first discuss the existence and cause of intrinsic stereotypes in LLMs, then move to the existing research on the fairness issues in LLM-based recommendations.
II-A Stereotype of LLMs
Stereotypes refer to social bias and discrimination that associate diverse combinations of characteristics with specific groups [8], such as a combination of “female nurse” or “male doctor”. Existing studies have revealed that LLMs trained on large-scale uncurated data, particularly uncensored and imbalanced text corpora, inherit various stereotypes against specific social groups, which can be categorized by age, gender, and religion, etc. [6, 7]. More specifically, stereotypes are ubiquitously encapsulated into the LLM vocabulary of word embeddings, leading to stereotypical behavior in various LLM-based applications since the fixed vocabulary is used to represent downstream tasks as textual information (e.g., task-specific prompts) [26].
Compared to conventional recommendation models, LLMs introduce unique stereotypes for inferring user preferences, leading to substantial discrimination in personalized recommendations. In other words, discriminative predictions can be elicited by prompting LLMs with textual information of users and items. For example, a female stereotype can be probed by LLMs given an interacted item “dress” in e-commerce recommendations. Notably, recent studies have indicated that most existing LLMs, such as T5, GPT, and LLaMA families, suffer from discriminative behaviors against stereotypes (e.g., gender) in recommendation tasks [11, 12, 21].
II-B Fairness in LLM-based Recommendation
Compared to unique user/item representations in conventional recommender systems, users and items share the textual representation (e.g., word embeddings) in LLM-based recommendations. Recent studies have revealed that LLMs are likely to inherit stereotypes that are embedded ubiquitously in word embeddings, due to their training on large-scale uncurated datasets [6, 7]. For example, LLM-RS are likely to stereotypically suggest “female nurse” and “male doctor” in job recommendations [7, 8]. This leads to LLM-RS exhibiting stereotypical linguistic associations between users and items, causing a form of two-sided (i.e., user-to-item) recommendation fairness. However, most existing studies on LLM-RS fairness are limited to either user-side or item-side fairness. These technical solutions involve augmenting imbalanced datasets or fine-tuning models with fair objectives, aiming to achieve consistent recommendation performance across user groups [12, 13] or fair exposure opportunities across item groups [14, 15]. Therefore, current fairness methods might fall short in addressing two-sided recommendation fairness, since LLM stereotypes can simultaneously involve user and item groups.
In addition to the aforementioned limitation on single-sided fairness, some existing works merely leverage discrete IDs or ID embeddings of users and items for recommendations [11, 16], which implicitly ignore stereotypes in textual information, such as user profiles and item descriptions. Notably, recent studies have indicated that textual knowledge is critical for harnessing the linguistic capabilities of LLM-RS to effectively comprehend user preferences in recommendations [27, 28]. To sum up, effective methods to tackle the recommendation fairness against intrinsic stereotypes of LLMs, specifically against biased textual knowledge, remain underexplored.
III PRELIMINARY
In this section, we elaborate on the fairness definitions for quantifying LLM stereotypes in recommendation tasks.
III-A Fairness Definition
There exist various definitions of recommendation fairness, such as consistent, calibrated, and counterfactual fairness, where the choice of these definitions depends on the scope of applications [9, 10]. For instance, in job recommendations, consistent fairness focuses on whether like cases are treated alike, such as whites potentially receiving more recommendations than blacks despite having equal qualifications. In this paper, we investigate the unique characteristics of LLM stereotypes, causing a form of two-sided (i.e., user-to-item) recommendation fairness.
To address fair recommendations, as illustrated in Fig. 2, we argue that user-to-item stereotypes should not amplify a user’s preference for items. Therefore, we extend the concept of calibrated fairness, requiring a calibrated proportion of stereotype groups between the original distribution (i.e., user-item interactions) and recommendation results. Formally, given any stereotype group $G$, stereotype-aware fairness requires:
$\mathbb{P}(\mathcal{V}_{u}^{\bm{\mathrm{interact}}}\in G)=\mathbb{P}(\mathcal{V}_{u}^{\bm{\mathrm{recommend}}}\in G)\quad\forall u\in\mathcal{U},$ (1)
where $\mathcal{V}_{u}^{\bm{\mathrm{interact}}}$ and $\mathcal{V}_{u}^{\bm{\mathrm{recommend}}}$ are the item sets in a user $u$’s historical interactions and recommendation results, respectively. $\mathcal{U}$ denotes the user set and $\mathbb{P}$ stands for the proportion.
III-B Recommendation Task
To adapt generative LLMs to recommendation tasks, recent advances have demonstrated the necessity to present target item candidates into prompts [29, 27] or additional tokens [28, 17] of LLM-RS. Therefore, we formulate recommendations as a binary classification task, where LLM-RS will determine whether or not to recommend a given target item $v$ based on the sequence of a user $u$’s historical interactions $\mathcal{H}_{u}$.
IV STEREOTYPE-AWARE FAIRNESS
In pursuit of fair recommendations, we extend the concept of calibrated fairness, requiring a calibrated proportion of stereotype groups between the original distribution (i.e., user-item interactions) and recommendation results.
IV-A Stereotype Group of User and Item
As shown in Fig. 1, LLM stereotypes can simultaneously involve user and item groups. For example, a “female” user and an item “dress” can be grouped into the same gender stereotype. Formally, let $G\in\mathcal{G}$ denote each stereotype group in a recommendation dataset, the user-side stereotype that $u\in G$ can be directly determined by user attributes, such as gender and age. As for item-side stereotype, we interpret it as a degree $d_{v\in G}$ to which this item is mostly interacted by users of a certain stereotype group. For example, in the case of binary stereotype groups $\mathcal{G}=[G_{1},G_{2}]$ where an item $v$ is interacted by $30\%$ users $u\in G_{1}$ and $10\%$ users $u\in G_{2}$, the degree $d_{v\in G_{1}}=0.2$ can be calculated by their subtraction. Therefore, for any stereotype group $G\in\mathcal{G}$ (where $G^{\prime}\neq G$), the degree to which an item $v$ is biased to a certain stereotype group $G$ can be calculated by
$d_{v\in G}=\mathop{\max}_{G}(\frac{\sum_{u\in G}{\mathbb{1}(v\in\mathcal{H}_{u})}}{\sum_{u\in G}1}-\sum_{G^{\prime}\in\mathcal{G}}{\frac{\sum_{u\in G^{\prime}}{\mathbb{1}(v\in\mathcal{H}_{u^{\prime}})}}{\sum_{u\in G^{\prime}}1}}),$ (2)
IV-B Stereotype Measurement
To measure the distribution of stereotype groups, given each interacted item $v$ in a user $u$’s interaction history $\mathcal{H}_{u}$, the proportion of any group $G$ can be formulated as:
$h_{u\in G}=\frac{1}{|\mathcal{H}_{u}|}\sum_{v\in\mathcal{H}_{u}}{\mathbb{1}(d_{v\in G})},$ (3)
where $\mathbb{1}$ represents an identity function
$\mathbb{1}(d_{v\in G})=\begin{cases}1,&\text{if}\ d_{v\in G}\geq\mathrm{threshold}\\ 0,&\text{otherwise}\end{cases}.$ (4)
In practice, a threshold should be applied to the above stereotype degree $d_{v\in G}$. This is because a small degree indicates a weak bias towards any stereotype group. For example, in Fig. 2, a “smartphone” is neutral to gender stereotypes.
Figure 3: Threshold of $d_{v\in G}$ based on Z-scores in different experimental datasets. An ablation study is conducted to validate the above settings of threshold in Fig. 6.
To determine the proper threshold, we employ the Z-score of $d_{v\in G}$ over all items in a recommendation dataset. Intuitively, Z-score identifies outliers based on how many standard deviations a data point is from the mean, where we regard outliers as items strongly related to certain stereotypes. As illustrated in Figure 3, we take the commonly-used settings $z=2$ and $z=3$ [30] to determine the threshold in Eq. (4) tailored to different experimental datasets.
TABLE I: Results of preliminary experiment. In the preliminary setup, LLM-RS are fine-tuned on recommendation datasets. The reported results are multiplied by 100, where boldface indicates the best score.
$G\in\mathcal{G}$ users $u$ target items $v$ implicit explicit counterfactual AUC Precision Recall
MovieLens $\mathcal{G}=[\mathrm{male},\mathrm{female}]$* $\in G$ $\notin G$ 12.02 17.14 10.76 46.93 60.08 64.84
$\in G$ $\in G$ 78.82 83.13 76.11 72.36 78.25 76.50
BookCrossing $\mathcal{G}=[\mathrm{teen},\mathrm{adult}]$* $\in G$ $\notin G$ 16.08 25.34 14.56 71.42 73.09 61.67
$\in G$ $\in G$ 63.10 67.15 69.19 89.80 95.11 70.86
  • * As detailed in Section VII-A, available gender and age information are provided by MovieLens and BookCrossing datasets, respectively.
IV-C Evaluation of Stereotype-aware Fairness
Based on the above stereotype measurement, extending the concept of calibrated fairness, we propose stereotype-aware fairness at the group level. In particular, given each user in any stereotype group $G$, the recommendation proportion of target items in the same group $G$ should be calibrated to the proportion of $\mathbf{G}$ (i.e., $h_{u\in G}$) in the user’s historical interactions. Formally, given the set of $N$ recommendations $\mathcal{S}=\{u_{i},v_{i}\}_{i=1}^{N}$ where each tuple denotes the target item $v$ (or $v_{i}$) recommended to a user $u$ (or $u_{i}$), the evaluation metric of stereotype-aware fairness is defined as:
$\bm{SF}=1-\frac{1}{|\mathcal{G}|}\sum_{G\in\mathcal{G}}{\frac{\sum_{u\in\mathcal{S}}{h_{u\in G}}}{\sum_{v\in\mathcal{S}}{\mathbb{1}(d_{v\in G})}}},$ (5)
where $\bm{SF}>0$ implies that LLM-RS amplify the recommendation proportion of any stereotype group $G$ compared to that (i.e., proportion of $G$) of user-item interactions, as marked by over-recommendation in Figure 2. Similarly, $\bm{SF}<0$ indicates under-recommendation.
V ANALYSIS OF STEREOTYPE-AWARE FAIRNESS
@@
In this section, we initiate a preliminary experiment to investigate the influence of stereotype-aware fairness, in addressing the following two research questions:
@@
  • RQ1: Does LLM-RS exhibit unfairness between the same ($u,v\in G$) and different ($u\in G,v\notin G$) stereotype groups?
@@
V-A Preliminary Experimental Setup
To validate the existence of the proposed stereotype-aware fairness in LLM-RS, exploratory experiments are conducted on two widely-used recommendation datasets: MovieLens and BookCrossing. The detailed description of datasets and evaluation metrics can be found in the main experiment section (please refer to Section VII-A).
V-B Analysis of RQ1
As illustrated in Table I, we compare the performance and fairness of LLM-RS when performing recommendations between users and items from different stereotype groups. As for performance comparisons, the recommendation quality between users and items from inconsistent stereotype groups is significantly inferior to that of consistent stereotype groups. In particular, the AUC, Precision, and Recall of LLM-RS decrease by 36%, 23%, and 15%, respectively, in recommendations between users $u\in G$ and target items $v\notin G$, such as recommending romantic movies (e.g., female stereotype) to male users. These observations imply that the recommendation quality of LLM-RS is sensitive to stereotypes, emphasizing the concern of stereotype-aware fairness toward the trustworthiness of LLM-based recommendations.
Despite the downgrade in performance, the fairness of LLM-RS regarding recommendations between inconsistent stereotype groups indicates a notably 62%-86% smaller value of $\bm{SF}$. In other words, the recommendation proportion of items from a stereotype group $G$ is much more calibrated to the proportion of $G$ in user-item interactions. However, negative $\bm{SF}$ can be observed in the MovieLens dataset, meaning that LLM-RS rarely recommend items $v\in G$ to users $u\notin G$ despite a large proportion of $G$ in the user’s historical interactions. The aforementioned differences in recommendations between consistent and inconsistent groups imply that LLM-RS tend to amplify the discrimination between user-side stereotypes and item-side stereotypes, such as exhibiting over-recommendation between users and items from the same stereotype group.
V-C Analysis of RQ2
Since stereotypes are simultaneously embedded in the word embeddings of users and items (e.g., user profiles and item titles), we aim to investigate the degree of stereotype-aware fairness under different levels of stereotypes, namely implicit, explicit, and counterfactual settings [21]. Specifically, the implicit setting only provides item titles in the input prompt of LLM-RS to infer the user-side stereotype without the actual user profile. In explicit and counterfactual settings, both user profiles and item titles are provided, as detailed in Section VII-A4. By comparing results between implicit and explicit settings, the unfairness of LLM-RS with implicit stereotypes decreases by 29%-36% for inconsistent stereotype groups and 5%-6% for consistent stereotype groups. These results support our findings that stereotypes exist in the word embeddings of both users and items, leading to stereotype-aware fairness in LLM-RS. Notably, LLM-RS shows a reduction in unfairness by a percentage of 2%-44% in a counterfactual world of explicit stereotypes, highlighting the potential of utilizing different stereotypes (e.g., counterfactual stereotype) to develop effective methods in addressing stereotype-aware fairness in LLM-RS.